Shift summary for Eve, 12/12/1999

Arun Tripathi (arun@physics.ucla.edu)
Sun, 12 Dec 1999 22:21:56 -0800 (PST)


Shift Summary for Eve shift, 12 Dec 99.

On Shift: BobH, Arun

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 7.5

Accounting of lost time: Waiting for TeVatron cool down and turning
on for beam as well as debugging SA3 trigger source and Etot.

========================================================================
Runs ended this shift: Run 15208 - KTeV_Ped, 5 spills. (for REG ADC)

Tapes brought to FCC this shift:
(including tapes for online split) none
========================================================================
Accesses during shift:

After we were locked up, Elizabeth and BoBH made accesses to put the
regenerator in, and also to check out the SA3 trigger source.
Sydney and Jason went down to adjust the ET power supply voltages.

========================================================================
Overview of shift:

No beam (well, maybe 3 half-spills)
Jaosn and Sydney spend several hours debugging CsI channels 788 and
1649. Elizabeth puts the reg. crate into readout, and turns on
the Reg HV.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15:30 Jaosn found bits 788 and 1649 missing from the laser
illumination plot.

Bit 788 had dead input buffer. Jason swapped the board #55
with spare board #30, and this bit is fixed.

Bit 1649 took several hours of Jason and Sydney's time to
fix. Turned out to be an irregular timing and low light
channel. Basically a false alarm, back to normal by
22:05.

16:10 Onsplit is modified to have the correct configuration
of tapedrives. Onsplit monitoring started.

JohnS stopped his B01 skim, and deleted all his files
from the online disks.

19:30 NM4AN unlocked by BoBT

19:35 Elizabeth restored HV to what remains of the regenerator.

22:30 Elizabeth comes in and puts the Reg crate (15) back into
the readout. REG ADC's appear in monitoring.

23:40 DC3 illumination looks funny, because SA3 wasn't counting.
Elizabeth and BoBH investigate. There was a problem with
a 622 coincidence module upstairs. SA3 trigger was moved to the
QUAD AND module previously used by ZCAL. SA3 looks normal
now.
23:50 Did PED run 15208.

Initialized BEAM_MUON run, waiting for BEAM.

========================================================================
Follow up Notes for next Shift


Do a PED run AFTER the muon run
Take the regenator OUT after the muon run.
Page ED after 0.5 hrs of HADRON beam.
Check the SA3 timing.
Load new DC alignment constants

========================================================================